International Data Systems' information on revenue and costs is relevant only up to a sales volume of 115,000 units. After 115,000 units, the market becomes saturated and the price per unit falls from $16.00 to $9.80. Also, there are cost overruns at a production volume of over 115,000 units, and variable cost per unit goes up from $8.00 to $8.25. Fixed costs remain the same at $65,000.

a. Compute operating income at 115,000 units.

Operating Income _________


b. Compute operating income at 215,000 units.

Operating Income __________
Answer #1

A. For units of 115,000 (unsaturated market): Revenue = Price per unit * No. of units = $16*115000 = $1,840,000

Fixed cost = $65000

Variable cost = variable cost per unit* No. of units = $8*115000 = $920,000

Operating income = Revenue - fixed cost - variable cost = $1,840,000 - $65,000 - $920,000 = $855,000

B. For units of 215,000 (Saturated market): Revenue = Price per unit * No. of units = $9.8*215000 = $2,107,000

Fixed cost = $65000

Variable cost = variable cost per unit* No. of units = $8.25*215000 = $1,773,750

Operating income = Revenue - fixed cost - variable cost = $2,107,000 - $65,000 - $1,773,750 = $268,250